INAF queue observing, 2018 November 08-09

Observer: R. Carini, M. Faccini
Telescope Operator: D. Gonzal-Huerta
Support Astronomer: B. Rothberg
Instrument: LUCI1 (LS mode), LUCI2

Summary

Night duration (12\xB0 el twilight) 1:15-12:45 UT = 11h 30m

Observing time: 10h 58min (94%)

Weather time loss: h m (0%)

technical time loss 0h 42m (6%)

Data Summary

At the sunset the sky is clear, wind between 5 - 15 m/s (front - rear) , humidity about 48%

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
calibration LUCI 1 & 2 DARK LUCI     luci1.20181109.0004-0028; luci2.20181109.0006-0030  
Polletta telluric HIP106817 luci2 mos _   luci2.00-00  
Polletta PLCK G73.4-57.5 luci2 mos G200 LoRes, zJspec 0.5/4 h mean seeng 0.6" COMPLETED
Cresci telluric HIP 6159 luci1 ls G200 LoRes, zJspec   luci1.0029-0031  
Cresci A1 QSO J0129 luci1 + luci2 G200 LoRes, zJspec 2h/2h luci1.0037-0044, luci1.0051-0071,luci2.0067-0087 completed
Cresci A1 ima QSO J0129 luci1 + luci2 J 800sec/800sec luci1.0072-0079,luci2.0088-0096 completed
Cresci telluric HIP6159 luci2 ls G200 LoRes, zJspec   luci2.0100-0102  
Pinna V577 Per luci2 AO N30, FeII, Ks, P-beta, H2 0.5/0.5 luci2.0104-0152 completed
Perna RCSGA327 luci2 ls G200 LoRes, zjspec 0.3/3h luci2.0157-0164  
Perna telluric HIP19053 lusi2 ls G200 LoRes, zjspec   luci2.0168-0171  
Severgnini psf UCAC 589-042207 luci2 AO N30, K 180s/180s luci2.0172-0174  
Severgnini A5 sdssj075223 luci2 AO N30, K 0.25/0.5h luci2.0179-0193  
Severgnini psf UCAC 589-042207 luci2 AO N30, K 180s/180s luci2.0194-196  
Nardini telluric HIP50 459 luci1 ls G150Ks, Ks   luci1.0083-0086  
Nardini q3   luci1 ls G150Ks, Ks 0.67/1.33h luci.0091-0105 completed
Luci flats and arc Pinna, Cresci, Polleta          

Details

Dark Luci
  • Perna luci1.20181108.0004-0008, luci2.20181108.0006-0010 dit 150 ndit 1 mer nor
  • Antoniucci D75b luci1.0009-0013,luci2.0011-0015 dit 50 ndit 1 mer nor
  • Antoniucci D75, Nardini q5==luci1.0014-0018,luci2.0016-0020== dit 100 ndit 1 mer nor
  • tell Perna luci1.0019-0023,luci2.0021-0025 dit 3 ndit 5 lir int
  • tell Nardini q3 luci1.0024-0028,luci2.0026-0030 dit 3 ndit 3 lir int

00:30 TO opens the dome
  • AO check
  • collimation

01:38 pointing check on the guide star of the telluric for Polletta's program

01:44 send the script for the telluric HIP106817
  • acq luci2.0031-0033
  • sci luci2.0034-0037
  • seeing 0.94" on the DIMM

02:00 slew to the target PLCK G73.4-57.5
  • luci2.0038-0040
  • 02:20 Field Stop Misalignment for N3.75 Imaging luci2.0041
  • 02:17 we load again the script
  • acq luci2.0042-0043== no good
  • acq luci2.0044-0048
  • sci luci2.0049-0063
  • seeng 0.67"on the DIMM
  • 02:57 seeing 0.60" on the dimm
  • 03:07 seeing about 0.65" on the dimm

03:25 slew to Cresci's program, telluric HIP 6159
  • TO authorizes both luci
  • TO de-rotates and collimates

03:38 send the script for the telluric
  • 03:41 mos error in LUCI2, we continue with luci1 only.
  • TO auhorizes luci1 only.

03:46 send the script again
  • acq luci1.0026-0028
  • sci luci1.0029-0031
  • seeing 0.88 on DIMM

04:00 sent the scientific script A1.
  • acq luci1.0032-0036
  • 04:14 sci luci1.0037-0044
  • seeing 0.7" on the DIMM

04:39 fixed mos error in luci2. we stop the integration. TO authorizes both luci

04:42 send again the script
  • acq luci1.0045-0050,luci2.0064-0066
  • 05:00 sci luci1.0051-0071,luci2.0067-0087
  • seeing 0.6" on DIMM
  • 05:16 seeing 0.7" on the DIMM
  • 05:22 seeing 0.6" on hte DIMM
  • 05:54 seeing 0.8" on the DIMM
  • 06:03 seeing 0.7" on the DIMM

06:11 send the script for the ima luci1.0072-0079,luci2.0088-0096

06:25 slew to telluric for luci2
  • TO authorizes LUCI2
  • 06:26 we send the script
  • acq luci2.0097-0099
  • sci luci2.0100-0102
  • seeing 0.6"

06:40 slew to AO mode. Pinna's program
  • AO check

06:46 send the script for the target V577 Per
  • sci luci2.103 test ima
  • 07:01 AO data luci2.0104-0138
  • seeing 1.0 on the DIMM, WFS 1.05",fwhm 0.73", estimated magnitude 8.4
  • DIMM 1.02", WFS 1.07", fwhm 0.87", estimaed magnitude 8.4
  • 07:15 dimm 1.1, WFS 1.1, fwhm 0.75"
  • 07:18 dimm 0.99" wfs 1.02" fwhm 0.71" estimated magnitude 8.6
  • 07:21 dimm 1.0" wfs 1.02" fwhm 0.86", estimated mag 8.5
  • loop opend, ima luci2.0118 is not good
  • 07:24 loop closed
  • dimm 0.91", wfs 0.91", fwhm 0.71"
  • 07:33 dimm 0.88", wfs 0.87", fwhm 0.69" estimated magnitude 8.5

07:45 error during commit, offset failed, loop opened

7:51 start again luci2.0139-0152

08:12 slew to the target of Perna
  • 8:16 due to long slewing, check point
  • 8:19 send the preset
  • acq==luci2.0153-0156==
  • sci luci2.0157-0164
  • 8:43 seeing 0.7"
  • 8:52 Seeing 0.7"
  • 8:59 seeing 0.6"

09:09 slew to telluric Perna HIP19053
  • acq luci2.0165-0167
  • sci luci2.0168-0171
  • seeing 0.7" on DIMM

09:21 slew to Severgnini AO program
  • pointing check

09:26 send the psf star for A5, UCAC 589-042207

*09:37 error in the preset, we send it again
  • 09:39 error again, we send again
  • 09:47, loop closes, start the integration
  • sci luci2.0172-0174
  • seeing 0.77 on the DIMM

09:50 alignment N30 ==luci2.0174-0177

10:08 slew to the scientific target
  • 10:10 start the science luci2.0179-0193 in the first image we see that the field stop of the camera is not alignment again.
  • 10:20 seeing 0.68" on dimm, fwhm 0.64", 0.7"
  • 10:23 offset failed , we pause
  • 10:23 we go on
  • 10:32 seeing 0.7"

10:43 trouble in the AO system.offset failed.
  • 10:49 send again the preset
  • lost the star, chck pointing.
  • 10:51 send again the preset
  • it is not possible close the loop because some cirrus arrived.

11:04 we can not close the loop due to cirrus
  • we try to go to the psf star, as required by the PI.
  • problem with the collimation
  • 11:29 colimation done
  • 11:23 loop close. luci2.0194-0196

11:29 slew to Nardini's program telluric for q3, HIP50 459, TO authorizes the LUCI1
  • preset failed, check pointing
  • 11:36 send again the preset
  • acq luci1.0079-0082
  • seein 0.9"-1"
  • 11:42 camera get lost
  • 11:45 problem fixed
  • sci luci1.0083-0086

11:50 slew to the target * acq luci1.0087-0090= * 11:59 sci luci.0091-0105 * seeing 0.5" * 12:18 seeing 0.75"

Calibrations

Pinna flat field P-beta N30 luci2.0197-0203 ATTENTION the keyword PI_NAME= Emanuele Nardini

Cresci flat ima, filter J luci1.0106-0115, luci2.0204-213
  • luci1.0116-0120, luci2.0214-218 flush
  • luci1.0121-0125, luci2.0219-223 flat zjspec, Halo1
  • luci1.0126-0127, luci2.0224-225 arc zjspec, Ne
  • luci1.0128-0129, luci2.0226-227 arc zjpec, Ar

Polletta we repeat the calibration because we have aligned for the N30 camera
  • luci2.0228-0232 flush
  • luci2.0233-0237 halo2 flat
  • luci2.0238-0239 arc Neon
  • luci2.0240-0241 arc Argon
Topic revision: r13 - 13 Dec 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback